# The Rees algebra and analytic spread of a divisorial filtration

### v1

@@ 3.1. Resolutions and Alterations
(Theorem 4.1 [22]) Let $X$ be a projective variety over a field $k$. Then there exists an alteration $\varphi:X^{\prime}\rightarrow X$ such that $X_{1}$ is a projective $k$-variety and a nonsingular scheme.
3.2. Proof of Theorem 1.4
In this subsection we finish the proof of Theorem 1.4. Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}$. Let $d=\dim R$.
Step 1. Suppose that $R$ is normal. By Lemma 3.3, there exists a birational projective morphism $Y_{1}\rightarrow\mbox{Spec}(R)$ such that $Y_{1}$ is normal and there exists an effective ${\mathbb{Q}}$-Weil divisor $D_{1}$ on $Y_{1}$ such that $\mathcal{I}=\mathcal{I}(D_{1})$.
The morphism $Y_{1}\rightarrow\mbox{Spec}(R)$ is the blow up of an ideal $J$ in $R$. $R$ is the local ring $\mathcal{O}_{V,q}$ of a (not necessarily closed) point $q$ on a projective $k$-variety $V$. Let $\mathcal{J}$ be an ideal sheaf on $V$ such that $\mathcal{J}_{q}=I$. Let $V_{1}\rightarrow V$ be the normalization of the blow up of $\mathcal{J}$. Let $V_{2}\rightarrow V_{1}$ be an alteration (whose existence follows from Theorem 3.5). This morphism is necessarily projective, since it is a morphism of projective varieties. Let $Z=V_{2}\times_{V}\mbox{Spec}(R)$. Then the induced morphism $Z\rightarrow Y_{1}$ is a projective alteration.
@@
Define the divisorial filtration $\mathcal{I}(S_{q})=\{I(S_{q})_{n}\}$ on the $d$-dimensional excellent normal local ring $S_{q}$ by the localization $I(S_{q})_{n}=I(S)_{n}\otimes_{S}S_{q}$. We have that $I(S_{q})_{m_{0}}:m_{S_{q}}^{\infty}\neq I(S_{q})_{m_{0}}$. The base change $Z_{q}=Z\times_{S}S_{q}\rightarrow\mbox{Spec}(S_{q})$ is a birational projective morphism such that $Z_{q}$ is nonsingular. Further, $\mathcal{I}(S_{q})=\mathcal{I}(D^{\prime}|Z_{q})$ for the ${\mathbb{Q}}$-Cartier divisor $D^{\prime}|Z_{q}$ on $Z_{q}$. By Theorem 3.1, $\ell(\mathcal{I}(S_{q}))=\dim S=d$. $S[\mathcal{I}(S_{q})]=S[\mathcal{I}(S)]\otimes_{S}S_{q}$ and $\dim S_{q}[\mathcal{I}(S_{q})]/q_{q}S_{q}[\mathcal{I}(S_{q})]=\ell(\mathcal{I}(S_{q}))=d$ implies there exists a chain of distinct prime ideals $P_{0}\subset P_{1}\subset\cdots\subset P_{d}$ in $S[\mathcal{I}(S)]$ such that $q\subset P_{0}$. From the inclusion $R[\mathcal{I}(R)]\subset S[\mathcal{I}(S)]$, we have a chain of prime ideals $Q_{0}=P_{0}\cap R[\mathcal{I}(R)]\subset Q_{1}=P_{1}\cap R[\mathcal{I}(R)]\subset\cdots\subset Q_{d}=P_{d}\cap R[\mathcal{I}(R)]$. Now $R[\mathcal{I}(R)]\rightarrow S[\mathcal{I}(S)]$ is an integral extension by Proposition 2.4. Thus the prime ideals $Q_{0}\subset Q_{1}\subset\cdots\subset Q_{d}$ are all distinct by Theorem A.6 [3]. Since these ideals all contain $q\cap R=m_{R}$, we have that $\ell(\mathcal{I}(R))\geq d$. Since we have the upper bound $\ell(\mathcal{I}(R))\leq d$ of (2), we have that $\ell(\mathcal{I}(R))=d$. This completes the proof of Theorem 1.4 in the case that $R$ is normal.
Step 2. Now suppose that $R$ is a domain, which is not normal. There exist divisorial valuations $\nu_{1},\ldots,\nu_{t}$ of $R$ and $a_{1},\ldots,a_{t}\in{\mathbb{Q}}_{>0}$ such that =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=I(\nu_{1})_{a_{1}n}\cap\cdots\cap I(\nu_{t})_{a_{t}n}$ for $n\geq 0$, with $I(\nu_{i})_{m}=\{f\in R\mid\nu_{i}(f)\geq m\}$.
Let $S$ be the normalization of $R$ in the quotient field of $R$. Let $J(\nu_{i})_{m}=\{f\in S\mid\nu_{i}(f)\geq m\}$. For $n\in{\mathbb{N}}$, let
@@
so that $J_{n}\cap R=I_{n}$. Let $\mathcal{J}=\{J_{n}\}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $q_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. Now $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration of the $d$-dimensional normal local ring $S_{q}$, so $\ell(\mathcal{J}_{q})=d$ by Step 1. Let $0\neq x\in R$ be in the conductor of $S$ over $R$, so that $x\in\mbox{Ann}_{R}(S/R)$ is a nonzero divisor on $R$. Then by Proposition 3.2, we have that $\ell(\mathcal{I})=d$.
Step 3. Suppose that $R$ is reduced and $\mathcal{I}$ has an expression (16). Let $S=\oplus_{i=1}^{t}R/P_{i}$ where $P_{i}$, $1\leq i\leq t$ are the minimal primes of $R$. Let $\mathcal{J}=\{J_{n}\}$ where
$J_{n}=\oplus_{i=1}^{t}[\cap_{j}I(v_{ij})_{n\lambda_{ij}}].$
We have that $J_{n}\cap R=I_{n}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $q_{q}\in\mbox{Ass}_{S}(S/J_{m_{0}})$.
@@
at $P_{i}$, we see that $M_{P_{i}}=0$ for all $i$. Thus $A_{P_{i}}=R_{P_{i}}$ for all $i$ so that $A\not\subset P_{i}$ for all $i$, and so $A\not\subset\cup P_{i}$ by Proposition 1.11 i) [1]. Thus there exists $x\in A\setminus\cup P_{i}$. Since $R$ is reduced, $x$ is a nonzero divisor of $R$. Now $S_{q}\cong(R/P_{i})_{m_{R/P_{i}}}$ for some $i$, and $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration on the $d$-dimensional local domain $S_{q}$, so that $\ell(\mathcal{J}_{q})=d$ by Step 2. Thus $\ell(\mathcal{I})=d$ by Proposition 3.2.
Step 4. Now suppose that $R$ is an arbitrary equidimensional local ring. Let =$\mathcal{I}=\{I_{n}\}$ have an expression (16). Let $\tilde{R}=R/\cap_{i=1}^{r}P_{i}$, which is a reduced local ring, with natural surjection $R\rightarrow\tilde{R}$. Since $P_{i}$ is contained in the kernel of $\varphi_{i}$ for all $i$, we have that $\cap_{i=1}^{r}P_{i}$ is contained in $\varphi^{-1}(I_{v_{ij}})$ for all $i,j$, and so $\cap_{i=1}^{r}P_{i}$ is contained in $I_{n}$ for all $n$. Let $\alpha_{i}:\tilde{R}\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq r$. Let $\tilde{\mathcal{I}}$ be the ${\mathbb{Q}}$-divisorial filtration on $\tilde{R}$ defined by $\tilde{I}_{n}=\cap_{i=1}^{r}\cap_{i,j}\alpha_{i}^{-1}(I(v_{ij})_{n\lambda_{ij}})$. We have that $R/I_{n}\cong\tilde{R}/\tilde{I}_{n}$ for all $n$ and $m_{\tilde{R}}=m_{R}\tilde{R}$. Thus $m_{R}$ an associated prime of $R/I_{n}$ implies $m_{\tilde{R}}$ is an associated prime of $\tilde{R}/\tilde{I}_{n}$, so (~$\ell(\tilde{\mathcal{I}})=\dim\tilde{R}=\dim R$ by Step 3. Since there is a natural surjection
@@ 4.1. Boundedness of the difference function
There exists a field $\tilde{k}$ such that $\tilde{k}\subset R$ and $[R/m_{R}:\tilde{k}]<\infty.$ Let $\mathcal{I}(R)=\mathcal{I}$. The proof of Step 1 of Theorem 1.4 in Section 3.2 shows that there exists a finite field extension $L$ of the quotient field $K$ of $R$ such that if $S$ is the normalization of $R$ in $L$ then there exists a resolution of singularities $Z\rightarrow\mbox{Spec}(S)$ and an effective ${\mathbb{Q}}$-divisor $D^{\prime}$ on $Z$ such that $\mathcal{I}(S)=\mathcal{I}(D^{\prime})$ and $I(S)_{n}\cap R=I(R)_{n}=I_{n}$ for all $n$. Thus we have inclusions
@@ 4.3. Geometric preliminaries
In this section, we prove Lemma 4.5. By Kodaira vanishing,
@@
Since $((D_{n}+K_{T}+Z)\cdot Z)$ is bounded from below (as the $D_{n}$ are ample) we have that $h^{1}(Z,\mathcal{O}_{T}(D_{n}+K_{T}+Z)\otimes\mathcal{O}_{Z}))$ is bounded by Lemma 3.1 [10]. Thus using equation (24), there exists a constant $c$ such that
@@ 4.5. The example
We give the numeric expression of 38, which is computed using the formulas (36) and other formulas from [9]. This formula is not necessary to compute the asymptotic behavior of the first difference function $\lambda_{R}(I(nF))/I((n+1)F)$.
@@ 5. Symbolic powers with infinitely many Rees valuations
where the limit is over the schemes $\mathcal{O}_{mE}$. By Serre duality on an elliptic curve, $H^{1}(E,\mathcal{O}_{E}(D))=0$ if $D$ is a divisor on $E$ of positive degree. From induction on $m$ in the short exact sequences
@@
for all $n$. By the Riemann Roch theorem on $E$, $h^{0}(E,\mathcal{O}_{E}(n(q-p)+p))=1$ for all $n$. Thus there exists a unique point $q_{n}\in E$ such that $\mathcal{O}_{E}(n(q-p)+p)\cong\mathcal{O}_{E}(q_{n})$. Since $q-p$ has infinite order in the Jacobian of $E$, we have that the $q_{n}$ are distinct points and $q_{n}\neq q$ for all $n$. By (46) and (47), we have surjections
@@
Finally, we prove equation (44). Let $Q=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))$. $Q$ is a prime ideal in $\oplus_{n\geq 0}P^{(n)}=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta))$ (as explained before Remark 6.6 [10]) and since $E$ is the only exceptional curve of $\varphi:X\rightarrow\mbox{Spec}(R)$, $Q=\sqrt{m_{R}(\oplus_{n\geq 0}(\oplus_{n\geq 0}P^{(n)})}$ by Proposition 6.7 [10].

### v2

@@ 3.2. Associated primes and maximal analytic spread
Let $R$ be an equidimensional excellent local ring which has the property that if $P$ is a minimal prime of $R$ and $X\rightarrow\mbox{Spec}(R/P)$ is a birational projective morphism, with $X$ integral, then there exists a projective regular alteration $X^{\prime}\rightarrow X$. Let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
Proof.
Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}$. Let $d=\dim R$.
Step 1. Suppose that $R$ is normal. By Lemma 3.3, there exists a birational projective morphism $Y_{1}\rightarrow\mbox{Spec}(R)$ such that $Y_{1}$ is normal and there exists an effective ${\mathbb{Q}}$-Weil divisor $D_{1}$ on $Y_{1}$ such that $\mathcal{I}=\mathcal{I}(D_{1})$. By the assumptions of the theorem, there exists a projective regular alteration $Z\rightarrow Y$.
@@
Define the divisorial filtration $\mathcal{I}(S_{q})=\{I(S_{q})_{n}\}$ on the $d$-dimensional excellent normal local ring $S_{q}$ by the localization $I(S_{q})_{n}=I(S)_{n}\otimes_{S}S_{q}$. We have that $I(S_{q})_{m_{0}}:m_{S_{q}}^{\infty}\neq I(S_{q})_{m_{0}}$. The base change $Z_{q}=Z\times_{S}S_{q}\rightarrow\mbox{Spec}(S_{q})$ is a birational projective morphism such that $Z_{q}$ is nonsingular. Further, $\mathcal{I}(S_{q})=\mathcal{I}(D^{\prime}|Z_{q})$ for the ${\mathbb{Q}}$-Cartier divisor $D^{\prime}|Z_{q}$ on $Z_{q}$. By Theorem 3.1, $\ell(\mathcal{I}(S_{q}))=\dim S=d$. $S[\mathcal{I}(S_{q})]=S[\mathcal{I}(S)]\otimes_{S}S_{q}$ and $\dim S_{q}[\mathcal{I}(S_{q})]/q_{q}S_{q}[\mathcal{I}(S_{q})]=\ell(\mathcal{I}(S_{q}))=d$ implies there exists a chain of distinct prime ideals $P_{0}\subset P_{1}\subset\cdots\subset P_{d}$ in $S[\mathcal{I}(S)]$ such that $q\subset P_{0}$. From the inclusion $R[\mathcal{I}(R)]\subset S[\mathcal{I}(S)]$, we have a chain of prime ideals $Q_{0}=P_{0}\cap R[\mathcal{I}(R)]\subset Q_{1}=P_{1}\cap R[\mathcal{I}(R)]\subset\cdots\subset Q_{d}=P_{d}\cap R[\mathcal{I}(R)]$. Now $R[\mathcal{I}(R)]\rightarrow S[\mathcal{I}(S)]$ is an integral extension by Proposition 2.4. Thus the prime ideals $Q_{0}\subset Q_{1}\subset\cdots\subset Q_{d}$ are all distinct by [3, Theorem A.6]. Since these ideals all contain $q\cap R=m_{R}$, we have that $\ell(\mathcal{I}(R))\geq d$. Since we have the upper bound $\ell(\mathcal{I}(R))\leq d$ of (2), we have that $\ell(\mathcal{I}(R))=d$. This completes the proof of Theorem 3.5 in the case that $R$ is normal.
Step 2. Now suppose that $R$ is a domain, which is not normal. There exist divisorial valuations $\nu_{1},\ldots,\nu_{t}$ of $R$ and $a_{1},\ldots,a_{t}\in{\mathbb{Q}}_{>0}$ such that =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=I_{R}(\nu_{1})_{a_{1}n}\cap\cdots\cap I_{R}(\nu_{t})_{a_{t}n}$ for $n\geq 0$. Let $S$ be the normalization of $R$ in the quotient field of $R$. Let $J(\nu_{i})_{m}=\{f\in S\mid\nu_{i}(f)\geq m\}$. For $n\in{\mathbb{N}}$, let
@@
so that $J_{n}\cap R=I_{n}$. Let $\mathcal{J}=\{J_{n}\}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. Now $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration of the $d$-dimensional normal local ring $S_{q}$. A birational projective morphism $X\rightarrow\mbox{Spec}(S)$ induces a birational projective morphism $X\rightarrow\mbox{Spec}(R)$, so $\ell(\mathcal{J}_{q})=d$ by Step 1. Let $0\neq x\in R$ be in the conductor of $S$ over $R$, so that $x\in\mbox{Ann}_{R}(S/R)$ is a nonzero divisor on $R$. Then by Proposition 3.2, we have that $\ell(\mathcal{I})=d$.
Step 3. Suppose that $R$ is reduced and equidimensional and $\mathcal{I}$ has an expression (16). Let $S=\oplus_{i=1}^{t}R/P_{i}$ where $P_{i}$, $1\leq i\leq t$ are the minimal primes of $R$. Let $\mathcal{J}=\{J_{n}\}$ where
$J_{n}=\oplus_{i=1}^{t}[\cap_{j}I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}}].$
We have that $J_{n}\cap R=I_{n}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$.
@@
at $P_{i}$, we see that $M_{P_{i}}=0$ for all $i$. Thus $A_{P_{i}}=R_{P_{i}}$ for all $i$ so that $A\not\subset P_{i}$ for all $i$, and so $A\not\subset\cup P_{i}$ by prime avoidance. Thus there exists $x\in A\setminus\cup P_{i}$. Since $R$ is reduced, $x$ is a nonzero divisor of $R$. Now $S_{q}\cong(R/P_{i})_{m_{R/P_{i}}}$ for some $i$, and $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration on the $d$-dimensional local domain $S_{q}$, so that $\ell(\mathcal{J}_{q})=d$ by Step 2. Thus $\ell(\mathcal{I})=d$ by Proposition 3.2 (this is where the assumption that $R$ is equidimensional is used).
Step 4. Now suppose that $R$ is an arbitrary equidimensional local ring. Let =$\mathcal{I}=\{I_{n}\}$ have an expression (16). Let $\tilde{R}=R/\cap_{i=1}^{t}P_{i}$, which is a reduced local ring, with natural surjection $R\rightarrow\tilde{R}$. Since $P_{i}$ is contained in the kernel of $\varphi_{i}$ for all $i$, we have that $\cap_{i=1}^{t}P_{i}$ is contained in $\varphi^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{i,j}})$ for all $i,j$, and so $\cap_{i=1}^{t}P_{i}$ is contained in $I_{n}$ for all $n$. Let $\alpha_{i}:\tilde{R}\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq t$. Let $\tilde{\mathcal{I}}$ be the ${\mathbb{Q}}$-divisorial filtration on $\tilde{R}$ defined by $\tilde{I}_{n}=\cap_{i=1}^{t}\cap_{i,j}\alpha_{i}^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}})$. We have that $R/I_{n}\cong\tilde{R}/\tilde{I}_{n}$ for all $n$ and $m_{\tilde{R}}=m_{R}\tilde{R}$. Thus $m_{R}$ an associated prime of $R/I_{n}$ implies $m_{\tilde{R}}$ is an associated prime of $\tilde{R}/\tilde{I}_{n}$, so (~$\ell(\tilde{\mathcal{I}})=\dim\tilde{R}=\dim R$ by Step 3. Since there is a natural surjection
@@ 4.1. Boundedness of the difference function
There exists a field $\tilde{k}$ such that $\tilde{k}\subset R$ and $[R/m_{R}:\tilde{k}]<\infty.$ Let $\mathcal{I}(R)=\mathcal{I}$. The proof of Step 1 of Theorem 3.5 in Section 3.2 and [24, Theorem 4.1], shows that there exists a finite field extension $L$ of the quotient field $K$ of $R$ such that if $S$ is the normalization of $R$ in $L$ then there exists a resolution of singularities $Z\rightarrow\mbox{Spec}(S)$ and an effective ${\mathbb{Q}}$-divisor $D^{\prime}$ on $Z$ such that $\mathcal{I}(S)=\mathcal{I}(D^{\prime})$ and $I(S)_{n}\cap R=I(R)_{n}=I_{n}$ for all $n$. Thus we have inclusions
@@ 4.3. Geometric preliminaries
By Kodaira vanishing,
@@
Since $((D_{n}+K_{T}+Z)\cdot Z)$ is bounded from below (as the $D_{n}$ are ample) we have that $h^{1}(\mathcal{O}_{T}(D_{n}+K_{T}+Z)\otimes\mathcal{O}_{Z}))$ is bounded by Lemma 3.1 [11]. Thus using equation (28), there exists a constant $c$ such that
@@ 4.5. The example
We give the numeric expression of 42, which is computed using the formulas (40) and other formulas from [10]. This formula is not necessary to compute the asymptotic behavior of the first difference function $\lambda_{R}(I(nF))/I((n+1)F)$.
@@ 5. Symbolic powers with infinitely many Rees valuations
where the limit is over the schemes $mE$. By Serre duality on an elliptic curve, $H^{1}(E,\mathcal{O}_{E}(D))=0$ if $D$ is a divisor on $E$ of positive degree. From induction on $m$ in the short exact sequences
@@
for all $n$. By the Riemann Roch theorem on $E$, $h^{0}(\mathcal{O}_{E}(n(q-p)+p))=1$ for all $n$. Thus there exists a unique point $q_{n}\in E$ such that $\mathcal{O}_{E}(n(q-p)+p)\cong\mathcal{O}_{E}(q_{n})$. Since $q-p$ has infinite order in the Jacobian of $E$, we have that the $q_{n}$ are distinct points and $q_{n}\neq q$ for all $n$. By (50) and (51), we have surjections
@@
Finally, we prove equation (48). Let $Q=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))$. $Q$ is a prime ideal in $\oplus_{n\geq 0}P^{(n)}=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta))$ (as explained before Remark 6.6 [11]) and since $E$ is the only exceptional curve of $\varphi:X\rightarrow\mbox{Spec}(R)$, $Q=\sqrt{m_{R}(\oplus_{n\geq 0}P^{(n)})}$ by Proposition 6.7 [11].